OpenStudy (anonymous):

Bob and Ally are organizing a school dance party. The expenses for the party are $88, and they are selling the tickets to the party for $11 each. If profit is modeled as a function, what is the domain of this function?

OpenStudy (anonymous):

so we are trying to find the x value ?

OpenStudy (anonymous):

P(c) = 24c + 1 P(c) = c + 24 P(c) = 24c – 1 P(c) = c – 24

OpenStudy (anonymous):

these are the possible answers

OpenStudy (anonymous):

thats the function

OpenStudy (anonymous):

do you know how to do this

OpenStudy (anonymous):

P(c)= 11x -88 so give integer values to x and u get the domain {-77, -66, ...}